# Comparing Machine Learning Algorithms by Union-Free Generic Depth

### 20 Dec 2023 v1

@@ 6.1.2 Discussion on Related Methods
The ufg depth approach presented in this paper provides a depth measure for all possible partial orders given by the items BS, CART, EN, GBM, GLM, LASSO, RF and RIDGE. The depth function is therefore a description of the distribution of all possible posets based on the observed posets. Applying the ufg depth to the posets obtained from the data sets provided by the UCI repository, we get that the maximum depth value of the observed and all possible posets is $0.32$. The poset corresponding to the maximum depth value is given by the left side of Figure 2. The minimum depth value is zero since any poset that states that CART dominates GBM or RIDGE has to have a depth value of zero according to Theorem 8.
@@
The original Bradley-Terry model, see Bradley and Terry (1952), was defined for total orders only. It is based on pairwise comparisons and assumes that these pairwise comparisons are independent. There are several approaches to extending the Bradley-Terry model, e.g. Davidson (1970); P. V. Rao and L.L. Kupper (1967). We focus on the extension of the Bradley-Terry model developed by Davidson (1970). This model was originally developed to analyze pairwise comparison data, where the participants can indicate a preference as well as have no preference. Sinclair (1982) showed that this approach can be rewritten as a generalized linear model using the Poisson distribution with log link. Here, ties are considered to occur when there is no preference between two items. This extended Bradley-Terry model provides us with so-called worth parameters and a discrimination parameter that includes the no preferences of all comparisons. In our situation, the worth parameters model the latent performance structure of a classifier compared to another classifier. More precisely, if $\pi_{i}$ denotes the worth of classifier $i$ (with $\sum_{i}\pi_{i}=1$) and $\nu$ the discrimination parameter, then the model assumes that the probability that classifier $i$ is preferred over classifier $j$ is given by $\pi_{i}/(\pi_{i}+\pi_{j}+\nu\sqrt{\pi_{i}\pi_{j}})$. Therefore, the classifier with the highest worth parameter is assumed to dominate all other classifiers in a pairwise comparison. Since the worth parameters are values in $[0,1]$, it is unlikely that two classifiers will have exactly the same worth, and thus in most cases a total order results.
@@
Note that generally, one can not make definite statements. There are extreme situations where one method gives a very dense order as a result whereas the other method gives a very sparse poset: If the single performance measures are highly anti-correlated, then the method presented in this paper might lead to very sparse poset data and thus also to a very sparse order as the deepest data point. In contrast, the method of GSD only relies on the marginal distribution of the single performance measures over the data sets and not on the correlation structure. Therefore, this method can still lead to very dense orderings as a result. On the other hand, if one has a strongly correlated structure of the performance measures and if additionally, the obtained orders are very similar on the majority of data sets, then the data depth method from here would give a very dense partial order. If at the same time, a small minority of measures is very different (i.e., outlying), then the method of Jansen et al. (2023a) might lead to a very sparse result because it uses the cardinal scale of the performance measures.
@@ 6.2.1 Data Set
In our comparison we are interested in the performance of the following supervised learning methods: Random Forests (RF, implemented in the R-package ranger Wright and Ziegler (2017)), Decision Tree (CART, implemented via the rpart library Therneau et al. (2015)), Logistic regression (LR), L1-penalized logistic regression (LASSO, implemented through the glmnet library Friedman et al. (2021)), and k-nearest neighbors (KNN, through the kknn library Hechenbichler and Schliep (2004)). As stated in the OpenML experiment-documentation all methods are run with default settings of the corresponding libraries. Hence, our application analyzes the behavior of methods using default settings and does not necessarily extend to general statements about the performance of hyperparameter-tuned versions of the respective algorithms. The algorithms were chosen as a selection of widely used supervised learning methods that perform reasonably without much tuning, in contrast to methods such as neural networks or boosting, which require considerable tuning to perform well.
@@
Finally, we want to highlight a structural observation about the order dimension of posets, which applies to all posets ordering five items. The order dimension of a finite poset is defined as the minimum number of linear extensions of this poset whose intersection is equal to this poset. In the case of posets based on five items, the order dimension is always less than or equal to two. [^20] We established this fact by simply computing for each poset on $5$ items its order dimension, see Footnote 1 for the code. This implies that each pose can be constructed by an intersection of at most two total orders. Thus, theoretically, two different performance measures could be sufficient to obtain all possible posets. (This is indeed possible, consider for example two measures where one gives more weight to true positives and the other gives more weight to true negatives. Then any combination of two linear orders can be obtained.) In contrast, for the data sets given by the UCI repository with eight items, two performance measures are not sufficient to obtain every possible poset on eight items. For example, the crown $S_{4}^{0}$, see Trotter (1974), is a poset on eight items with order dimension four. Therefore, in the case of Section 6.1, at least four performance measures are required. Note that the above observation should not be taken as advice on how to choose the (number of) performance measures. One aspect is that, given a set of performance measures $A$, it is questionable whether there are two performance measures that represent exactly the same order as that obtained by the set of performance measures $A$. This issue of reducing the number of performance measures based on the order dimension is discussed in the next section. In any case, the decision on which and how many measures to include in an analysis should always be based on substance considerations.
@@ 6.2.2 Analysis using UFG Depth
Figure 6: OpenML: Each point in the scatter plot describes the depth of one poset based on the observed posets given by two different performance sets but based on the same data sets. The scatter plot on the right upper corner compares the depth of the posets based on the observed ones using area under the curve and F scorre with the observed posets using all four performance measures.
@@ Acknowledgments
We sincerely thank all four anonymous reviewers of the ISIPTA conference 2023 in Oviedo, Spain for valuable comments that helped to improve the paper. Moreover, we want to gratefully thank all participants at the ISIPTA’23 conference in Oviedo for all very helpful discussions. Hannah Blocher and Georg Schollmeyer gratefully acknowledge the financial and general support of the LMU Mentoring program. Hannah Blocher sincerely thanks Evangelisches Studienwerk Villigst e.V. for funding and supporting.

### 21 Feb 2024 v3

@@ 7.1.2 Discussion on Related Methods
The ufg depth approach presented in this paper provides a depth measure for all possible partial orders given by the items BS, CART, EN, GBM, GLM, LASSO, RF, and RIDGE. The depth function is therefore a description of the distribution of all possible posets based on the observed posets. Applying the ufg depth to the posets obtained from the data sets provided by the UCI repository, we get that the maximum depth value of the observed and all possible posets is $0.32$. The poset corresponding to the maximum depth value is given by the left side of Figure 2. The minimum depth value is zero since any poset that states that CART dominates GBM or RIDGE has to have a depth value of zero according to Theorem 9.
@@
The original Bradley-Terry model, see Bradley and Terry (1952), was defined for total orders only. It is based on pairwise comparisons and assumes that these pairwise comparisons are independent. There are several approaches to extending the Bradley-Terry model, e.g. Davidson (1970); Rao and Kupper (1967). We focus on the extension of the Bradley-Terry model developed by Davidson (1970). This model was originally developed to analyze pairwise comparison data, where the participants can indicate a preference as well as have no preference. Sinclair (1982) showed that this approach can be rewritten as a generalized linear model using the Poisson distribution with log link. Here, ties are considered to occur when there is no preference between two items. This extended Bradley-Terry model provides us with so-called worth parameters and a discrimination parameter that includes the no preferences of all comparisons. In our situation, the worth parameters model the latent performance structure of a classifier compared to another classifier. More precisely, if $\pi_{i}$ denotes the worth of classifier $i$ (with $\sum_{i}\pi_{i}=1$) and $\nu$ the discrimination parameter, then the model assumes that the probability that classifier $i$ is preferred over classifier $j$ is given by $\pi_{i}/(\pi_{i}+\pi_{j}+\nu\sqrt{\pi_{i}\pi_{j}})$. Therefore, the classifier with the highest worth parameter is assumed to dominate all other classifiers in a pairwise comparison. Since the worth parameters are values in $[0,1]$, it is unlikely that two classifiers will have exactly the same worth, and thus in most cases a total order results.
@@
Note that generally, one can not make definite statements. There are extreme situations where one method gives a very dense order as a result whereas the other method gives a very sparse poset: If the single performance measures are highly anti-correlated, then the method presented in this paper might lead to very sparse poset data and thus also to a very sparse order as the deepest data point. In contrast, the method of GSD only relies on the marginal distribution of the single performance vectors of the single classifiers over the data sets and not on the correlation structure between these vectors for different classifiers. Therefore, this method can still lead to very dense orderings as a result. On the other hand, if one has a strongly correlated structure of the performance measures and if additionally, the obtained orders are very similar on the majority of data sets, then the data depth method from here would give a very dense partial order. If at the same time, a small minority of measures is very different (i.e., outlying), then the method of Jansen et al. (2023a) might lead to a very sparse result because it uses the cardinal scale of the performance measures.
@@ 7.2.1 Data Set
In our comparison we are interested in the performance of the following supervised learning methods: Random Forests (RF, implemented in the R-package ranger Wright and Ziegler (2017)), Decision Tree (CART, implemented via the rpart library Therneau et al. (2015)), Logistic regression (LR), L1-penalized logistic regression (LASSO, implemented through the glmnet library Friedman et al. (2021)), and k-nearest neighbors (KNN, through the kknn library Hechenbichler and Schliep (2004)). For all methods, the choice of parameter settings depends on the goal of the user who uploaded the experiment and its results. Since different users may have different goals, this analysis does not necessarily extend to general statements about the performance of hyperparameter-tuned versions of the algorithms, where the goal is to increase their performance. The algorithms were chosen as a selection of widely used supervised learning methods that perform reasonably without much tuning, in contrast to methods such as neural networks or boosting, which require considerable tuning to perform well.
@@
Finally, we want to highlight a structural observation about the order dimension of posets, which applies to all posets ordering five items. The order dimension of a finite poset is defined as the minimum number of linear extensions of this poset whose intersection is equal to this poset. In the case of posets based on five items, the order dimension is always less than or equal to two. [^20] We established this fact by simply computing for each poset on $5$ items its order dimension, see Footnote 5 for the code. This implies that each poset can be constructed by an intersection of at most two total orders. Thus, theoretically, two different performance measures could be sufficient to obtain all possible posets. (This is indeed possible, consider for example two measures where one gives more weight to true positives and the other gives more weight to true negatives. Then any combination of two linear orders can be obtained.) In contrast, for the data sets given by the UCI repository with eight items, two performance measures are not sufficient to obtain every possible poset on eight items. For example, the crown $S_{4}^{0}$, see Trotter (1974), is a poset on eight items with order dimension four. Therefore, in the case of Section 7.1, at least four performance measures are required. Note that the above observation should not be taken as advice on how to choose the (number of) performance measures. One aspect is that, given a set of performance measures $A$, it is questionable whether there are two performance measures that represent exactly the same order as that obtained by the set of performance measures $A$. This issue of reducing the number of performance measures based on the order dimension is discussed in the next section. In any case, the decision on which and how many measures to include in an analysis should always be based on substance considerations.
@@ 7.2.2 Analysis using UFG Depth
Figure 6: OpenML: Each point in the scatter plot describes the depth of one poset based on the observed posets given by two different performance sets but based on the same data sets. The scatter plot on the right upper corner compares the depth of the posets based on the observed ones using area under the curve and F score with the observed posets using all four performance measures.
@@ Acknowledgments
We sincerely thank all four anonymous reviewers of the ISIPTA conference 2023 in Oviedo, Spain for valuable comments that helped to improve the paper. Moreover, we want to thank all participants at the ISIPTA’23 conference in Oviedo for all very helpful discussions. We are also grateful to the two reviewers for their review of this extended version. Hannah Blocher and Georg Schollmeyer gratefully acknowledge the financial and general support of the LMU Mentoring program. Hannah Blocher sincerely thanks Evangelisches Studienwerk Villigst e.V. for funding and supporting.
@@
    Writing - Review & Editing: Hannah Blocher (equal); Georg Schollmeyer (equal); Christoph Jansen (equal)
  • 9.